# Topology-Preserving Image Segmentation with Spatial-Aware Persistent Feature Matching

### v2

@@ Abstract
Topological correctness is critical for segmentation of tubular structures. Existing topological segmentation loss functions are primarily based on the persistent homology of the image. They match the persistent features from the segmentation with the persistent features from the ground truth and minimize the difference between them. However, these methods suffer from an ambiguous matching problem since the matching only relies on the information in the topological space. In this work, we propose an effective and efficient Spatial-Aware Topological Loss Function that further leverages the information in the original spatial domain of the image to assist the matching of persistent features. Extensive experiments on images of various types of tubular structures show that the proposed method has superior performance in improving the topological accuracy of the segmentation compared with state-of-the-art methods [^1] Our codes will be available at the time of publication..
@@ 1 Introduction
The primary research direction in topology-aware segmentation is the loss function [24, 28, 13, 14, 12, 34, 26, 33, 2, 3, 30]. Typically, a topological constraint is used in addition to the pixel classification loss. The topological constraint can be either a direct measure [13, 12, 34, 26, 2, 3, 30] of the topology of the object or an indirect approximation [24, 28, 33]. The direct methods mostly rely on the persistent homology [8] of the images and attempt to match and minimize the difference between the persistent features computed from the ground truth and the segmentation. A brief introduction of persistent homology will be provided in Sec. 3.1.2. Although many matching algorithms were proposed [13, 12, 34, 26, 2], they only rely on the global information in the topological space, i.e., the lifespan of the persistent features, which is ambiguous since the binary ground truth masks yield overlapping persistent features. This could cause more incorrect topological features to be prolonged and correct topological features to be eliminated, making the topological loss function less effective.
@@
Figure 1: Overview of our method. ‘Dgm’ is the persistent diagram of the images and the dotted lines indicate the spatial correspondences.
@@ 2 Related Work
For indirect methods, an introductory work is [24], which proposed to use an ImageNet pretrained VGG [6, 18] network to extract the topological information of the segmentation and ground truth. It is shown that different layers of the extraction network incorporate perceptual features which are approximations of the image topological features. The loss is then computed between the perceptual features of the prediction and the ground truth. [33] proposed a topological loss more exclusive for gland segmentation. The loss is computed based on the iterations needed to derive the skeleton of the glands by the erosion operation. [28] proposed a clDice loss which approximates the topological features of tubular structures by a relationship between the segmentation and its soft skeleton. It is shown that such a relationship helps strengthen the connectivity within tubular structures, resulting in better topological accuracy in the segmentation.
@@ 3.1.1 Topology in Digital Images
Figure 3: Visualization of persistent features matching. A matched 1-feature is marked with the same color with the feature in GT.
@@ 3.1.2 Persistent Homology
This sequence of complexes is called the filtration of $f$, or to be more specific, the sub-level filtration of $f$. Alternatively, a super-level filtration is the sequence of sub-complexes considering an $f$ with $f(c_{i})\geq f(c_{j})$ and the super-level sets $C(a)=f^{-1}\left[a,\inf\right)$. The value $a$ is said to be the filtration value. An easier way to look at it is that, given an image with discrete pixel values in $\left[0,1\right]$, we threshold the image from 1 to 0. The 0-cubes (pixels) and their cofaces larger than the threshold value are included in the cubical sub-complex. As the threshold value decreases, more cubes are included to form the current sub-complex. Finally, at the smallest threshold value, all cubes from the image are included and the final sub-complex becomes the full cubical complex (Fig. 2(b)). A homology feature can be born at a filtration value when a new cube, called the creator, is included into the current sub-complex, and it can be killed at a future filtration value when this feature is destroyed by the inclusion of another higher-dimensional cube, called the destroyer, into a future sub-complex. For example, a pixel whose value is a local maxima gives birth to a 0-homology feature (connected component) at a certain threshold value. As the value decreases, more and more surrounding pixels are included until sometime this 0-feature is connected to another 0-feature in the sub-complex by the introduction of an edge (1-cube) ‘filled’ between two pixels that connects the two areas. Then one of the features is killed (conventionally the newer one). The persistent homology is a method that describes this process where all the homology-features in the image are born and killed in the filtration. A homology-feature that lives in the filtration is said to be a persistent feature and the lifespan of that persistence feature ($born\ filtration\ value-kill\ filtration\ value$) is said to be the persistence of that feature.
@@ 3.1.3 Persistent Barcodes and Diagrams
Persistent barcodes ($\mathcal{B}$) are useful tools to describe and visualize the persistent homology of an image. As shown in Fig. 2(c), each barcode represents a persistent feature of the image, where the left edge corresponds to its birth filtration value and the right edge to its death filtration value. The persistent barcodes can be further transformed into a persistent diagram ($\mathcal{D}$, Fig. 2(d)). In $\mathcal{D}(\mathcal{I})$, each persistent feature in the image is represented by a dot whose coordinate is $(1-Birth\ value,1-Death\ value)$.
@@ 3.2 Spatial-Aware Topological Loss Function
Recent topology-aware segmentation methods generally follow the idea of matching the persistent diagrams or barcodes between the target and prediction and try to minimize the difference between them. However, (except for BMLoss) the matching is solely based on the distances between persistent features within the diagram or barcodes, i.e., the birth, death values and the persistence of the feature. Given the segmentation model output-likelihood L, the binarized prediction P and the ground truth T, these methods can make P to have more similar Betti numbers with the T. However, there is no guarantee that each persistent feature from L is matched to the correct one from T. This issue is prominent in grid-like data. For example, in a 2D image segmentation task, the ground truth mask is usually a binary image, which leads all points in $\mathcal{D}(T)$ to overlap at the upper-left corner of the diagram, i.e., (0, 1). As a result, the distances between each point in $\mathcal{D}(L)$ and any point in $\mathcal{D}(T)$ are the same, which cause confusions when matching the diagrams. We therefore propose to reduce this problem by introducing spatial awareness that further leverages the spatial distance between two persistent features in the original image when matching persistent diagrams.
@@ 3.2.1 Spatial-Aware Matching and Loss Formulation
Fig. 3 shows matching examples between T and L from a converged segmentation model (more detailed examples in supplementary material (SM)). From the limited examples we manually compared, it seems the Betti-matching [30] is indeed a more accurate method. But our proposed method still significantly reduces the matching error compared with the Wasserstein matching at a much smaller computational cost than the Betti-matching. Unfortunately, there is no ‘ground truth’ available to compare the matching accuracy numerically. Moreover, it is worth mentioning that although the matching on segmentation from a converged model offers a partial insight on this problem, it does not provide a comprehensive understanding of how the matching precisely influences the whole training process, which can be less straightforward to see. Accordingly, the final segmentation quality on testing images serves as the gold standard for the evaluation and empirically it is shown that a less confused matching enables better topological accuracy in the segmentation ([30], Table 1).
@@ 3.2.2 Gradient of SATLoss
where $\omega$ is the model weights, $L(c_{b}(\cdot)),L(c_{d}(\cdot))$ are the pixel values of the creator and destroyer in L. In practice, the gradient pushes the value of creator pixel up and destroyer down if the persistent feature has a match in $\mathcal{D}(T)$ to prolong the persistence of the persistent feature. On contrary, if the persistent feature is matched to the diagonal, the gradient pushes the creator value down and the destroyer value up to shorten the persistence of this feature. The difference with Wasserstein matching is that our SATLoss enables the gradient to act more selectively on the persistent features and their critical cells, reducing the number of persistent features that are incorrectly prolonged or eliminated.
@@ 4.1 Datasets
Our main comparison with the SOTA topological segmentation loss functions excluding the BMLoss is performed over four datasets with diverse tubular structures:
@@
  • CREMI [9]: Segmentation of cell boundaries of brain neurons in electronic microscopy images.
@@
Considering that the computational cost of persistent homology is still notably higher than conventional segmentation training objectives, it is a common practice in the literature of this field to crop the full image to small image patches for training and computing the evaluation metrics, for example, $64\times 64$ [28, 13]. However, too small patch size leads to limited semantic and topological information. In our experiments, we push it to the limit of our computational resources and use a patch size of around $300\times 300$ among training, validation and testing.
@@
$\mathcal{L}_{topo}$ Acc.$\uparrow$ Dice(P)$\uparrow$ clDice(P)$\uparrow$ $\beta_{0}$(P)$\downarrow$ $\beta_{1}$(P)$\downarrow$ Dice(F)$\uparrow$ clDice(F)$\uparrow$ $\beta_{0}$(F)$\downarrow$ $\beta_{1}$(F)$\downarrow$
TCLoss 96.10$\pm$0.64 48.66$\pm$5.06 63.66$\pm$6.93 2.591$\pm$0.87 0.487$\pm$0.31 63.73$\pm$5.51 68.55$\pm$5.98 197.7$\pm$63.0 54.7$\pm$27.1
clDice Loss 95.50$\pm$0.52 46.71$\pm$3.04 66.04$\pm$3.11 1.941$\pm$0.94 0.488$\pm$0.30 61.04$\pm$3.74 68.88$\pm$4.15 179.4$\pm$85.2 52.3$\pm$27.3
BMLoss 94.91$\pm$0.87 46.95$\pm$3.40 64.93$\pm$4.83 0.896$\pm$0.31 0.520$\pm$0.28 60.93$\pm$3.69 69.10$\pm$3.65 61.5$\pm$23.1 40.8$\pm$23.0
SATLoss (ours) 96.12$\pm$0.53 47.73$\pm$3.91 66.17$\pm$3.16 0.815$\pm$0.13 0.481$\pm$0.31 63.34$\pm$4.09 68.95$\pm$3.95 88.9$\pm$17.9 51.9$\pm$25.6
Table 3: Comparison with BMLoss on C.Elegan-small dataset.
$\mathcal{L}_{topo}$ Acc.$\uparrow$ Dice(P)$\uparrow$ clDice(P)$\uparrow$ $\beta_{0}$(P)$\downarrow$ $\beta_{1}$(P)$\downarrow$ Dice(F)$\uparrow$ clDice(F)$\uparrow$ $\beta_{0}$(F)$\downarrow$ $\beta_{1}$(F)$\downarrow$
TCLoss 98.66$\pm$0.04 90.95$\pm$1.63 94.64$\pm$0.84 0.470$\pm$0.05 0.264$\pm$0.09 92.65$\pm$0.52 96.14$\pm$0.50 1.517$\pm$0.35 0.906$\pm$0.25
clDice Loss 98.59$\pm$0.04 90.22$\pm$1.29 94.59$\pm$0.83 0.481$\pm$0.06 0.252$\pm$0.06 92.12$\pm$0.53 95.92$\pm$0.41 1.500$\pm$0.06 0.922$\pm$0.10
BMLoss 98.24$\pm$0.11 88.03$\pm$1.28 92.16$\pm$0.86 0.627$\pm$0.20 0.256$\pm$0.10 90.17$\pm$0.78 94.05$\pm$0.85 2.072$\pm$0.92 0.917$\pm$0.29
SATLoss (ours) 98.67$\pm$0.05 90.47$\pm$1.94 94.97$\pm$0.97 0.435$\pm$0.14 0.235$\pm$0.08 92.39$\pm$0.60 96.25$\pm$0.31 0.839$\pm$0.19 0.856$\pm$0.23
@@ 4.2 Implementation Details
Since the topic is the loss function, we follow [28, 13, 30, 12, 14] and adopt a consistent, general U-Net [27] network structure and set of hyperparameters for all experiments. An Adam optimizer is used with a weight decay of 1e-3 and an initial learning rate of 1e-3, decayed to 1e-4 at half of the total number of epochs. We train 30 epochs for Roads, 50 epochs for CREMI and CrackTree and 100 epochs for DRIVE. We use the best pixel loss reported in the corresponding paper for each of the baseline methods and use the BCE loss as pixel loss with our proposed method. The best weights $\lambda$ for the topological loss functions are determined by the ablation studies in Sec. 4.6. All experiments are repeated with 3 random seeds. For the smaller CrackTree and DRIVE datasets, we further adopt a 3-fold cross validation under each seed. In each independent random experiment, the model with the best Dice score on the validation set is used for testing and the average test scores and their stds over these experiments are reported. Our implementation is under the PyTorch framework, where the computation of persistent homology relies on the GUDHI (Geometry Understanding in Higher Dimensions) library [25].
For Roads-small and C.Elegan-small datasets, we follow [30] to determine the network structures and training hyperparameters. We train all loss functions using the same hyperparameters reported in [30] without finetuning in our method. Additionally, in [30] experiments were run for only once. Consider the small dataset size, we instead use a 3-fold cross validation where each fold is repeated with 3 random seeds to make the results statistically more convincing.
@@ 4.3 Evaluation Metrics
Two types of metrics: volumetric and topological are used to evaluate the general segmentation quality and the topological accuracy, respectively. For volumetric metrics, we adopt the pixel-wise accuracy and the Dice score. We also use a clDice score [28], which although is a pixel-based metric, indirectly implies the topological correctness. For topological metrics, we use the 0-th and 1-st Betti number error ($\beta_{0}$, $\beta_{1}$), which are the most commonly-used measurements of topological accuracy. We quit using the Betti-matching metric proposed in [30] because the computational cost is too high given the amount of our experiments.
@@ 4.5 Comparison with BMLoss
For comparison with BMLoss, we also include the clDice and TCLoss in the baseline, as the first method is the newest baseline in [30] and TCLoss is a newer method than BMLoss (which did not include BMLoss in their baseline). On another note, although the metrics are commonly computed on the image patches [13, 26, 28], in [30] they are computed over the full image. We therefore report the results on both the patch size P and full image size F. The result are shown in Table 2, Table 3 and Fig. 5. While our proposed method does not achieve the same performance level in global topological accuracy on full image size with BMLoss on the Roads-small dataset, on boarder evaluations which take the other results into consideration, our method demonstrates comparable, if not superior, performance. Furthermore, as reported in Table 4, our method achieves the results at a much smaller training time compared with BMLoss. Such a fundamental constraint of BMLoss is much less a problem of our method. It is important to note that in addition to the algorithm complexity, the coding language and implementation details can also influence the training time. The bottom-level codes in the GUDHI library we use to compute persistent homology are written in C++, while the codes for BMLoss are all written in Python. So this comparison is unfair in a strict sense. Given the complexity of the BMLoss, a reimplementation in C++ is challenging and is beyond the scope of this study. To the best of our knowledge, the Python version we used is the only and official implementation available. Therefore, despite the differences in implementation context, our proposed method still demonstrates superior computational efficiency.

### v3

@@ Abstract
Topological correctness is critical for segmentation of tubular structures, which pervade in biomedical images. Existing topological segmentation loss functions are primarily based on the persistent homology of the image. They match the persistent features from the segmentation with the persistent features from the ground truth and minimize the difference between them. However, these methods suffer from an ambiguous matching problem since the matching only relies on the information in the topological space. In this work, we propose an effective and efficient Spatial-Aware Topological Loss Function that further leverages the information in the original spatial domain of the image to assist the matching of persistent features. Extensive experiments on images of various types of tubular structures show that the proposed method has superior performance in improving the topological accuracy of the segmentation compared with state-of-the-art methods. Code is available at https://github.com/JRC-VPLab/SATLoss.
@@ 1 Introduction
The primary research direction in topology-aware segmentation is the loss function [25, 29, 13, 14, 12, 35, 27, 34, 2, 3, 31, 23]. Typically, a topological constraint is used in addition to the pixel classification loss. The topological constraint can be either a direct measure [13, 12, 35, 27, 2, 3, 31] of the topology of the object or an indirect approximation [25, 29, 34]. The direct methods mostly rely on the persistent homology [8] of the images and attempt to match and minimize the difference between the persistent features computed from the ground truth and the segmentation. A brief introduction of persistent homology will be provided in Sec. 3.1.2. Although many matching algorithms were proposed [13, 12, 35, 27, 2], they only rely on the global information in the topological space, i.e., the lifespan of the persistent features, which is ambiguous since the binary ground truth masks yield overlapping persistent features. This could cause more incorrect topological features to be prolonged and correct topological features to be eliminated, making the topological loss function less effective.
@@
Figure 1: Overview of our method. ‘Dgm’ is the persistent diagram of the images and the dotted lines indicate the spatial correspondences. Colored triangles indicate creators of topological features and are used as spatial location references for topological feature matching.
@@ 2 Related Work
For indirect methods, an introductory work [25] proposed to use an ImageNet pretrained VGG [6, 18] network to extract the topological information of the segmentation and ground truth. It is shown that different layers of the extraction network incorporate perceptual features which are approximations of the image topological features. The loss is then computed between the perceptual features of the prediction and the ground truth. [34] proposed a topological loss more exclusive for gland segmentation. The loss is computed based on the iterations needed to derive the skeleton of the glands by the erosion operation. [29] proposed a clDice loss which approximates the topological features of tubular structures by a relationship between the segmentation and its soft skeleton. It is shown that such a relationship helps strengthen the connectivity within tubular structures, resulting in better topological accuracy in the segmentation.
@@ 3.1.1 Topology in Digital Images
Figure 3: Visualization of persistent features matching. A matched 0/1-feature is marked with the same color with the feature in GT.
@@ 3.1.2 Persistent Homology
This sequence of complexes is called the filtration of $f$, or to be more specific, the sub-level filtration of $f$. Alternatively, a super-level filtration is the sequence of sub-complexes considering an $f$ with $f(c_{i})\geq f(c_{j})$ and the super-level sets $C(a)=f^{-1}\left[a,\inf\right)$. The value $a$ is said to be the filtration value. An easier way to look at it is that, given an image with discrete pixel values in $\left[0,1\right]$, we threshold the image from 1 to 0. The 0-cubes (pixels) and their cofaces larger than the threshold value are included in the cubical sub-complex. As the threshold value decreases, more cubes are included to form the current sub-complex. Finally, at the smallest threshold value, all cubes from the image are included and the final sub-complex becomes the full cubical complex (Fig. 2(b)). A homology feature can be born at a filtration value when a new cube, called the creator, is included into the current sub-complex, and it can be killed at a future filtration value when this feature is destroyed by the inclusion of another higher-dimensional cube, called the destroyer, into a future sub-complex. For example, a pixel whose value is a local maxima gives birth to a 0-homology feature (connected component) at a certain threshold value. As the value decreases, more and more surrounding pixels are included until sometime this 0-feature is connected to another 0-feature in the sub-complex by the introduction of an edge (1-cube) ‘filled’ between two pixels that connects the two areas. Then one of the features is killed (conventionally the newer one). The persistent homology is a method that describes this process where all the homology-features in the image are born and killed in the filtration. A homology-feature that lives in the filtration is said to be a persistent feature and the lifespan of that persistence feature (bornfiltrationvaluekillfiltrationvalueborn\ filtration\ value-kill\ filtration\ value) is said to be the persistence of that feature.
@@ 3.1.3 Persistent Barcodes and Diagrams
Persistent barcodes ($\mathcal{B}$) are useful tools to describe and visualize the persistent homology of an image. As shown in Fig. 2(c), each barcode represents a persistent feature of the image, where the left edge corresponds to its birth filtration value and the right edge to its death filtration value. The persistent barcodes can be further transformed into a persistent diagram ($\mathcal{D}$, Fig. 2(d)). In $\mathcal{D}(\mathcal{I})$, each persistent feature in the image is represented by a dot whose coordinate is (1Birthvalue,1Deathvalue)(1-Birth\ value,1-Death\ value).
@@ 3.2 Spatial-Aware Topological Loss Function
Recent topology-aware segmentation methods generally follow the idea of matching the persistent diagrams or barcodes between the target and prediction and try to minimize the difference between them. However, (except for BMLoss) the matching is solely based on the distances between persistent features within the diagram or barcodes, i.e., the birth, death values and the persistence of the feature. Given the segmentation model output-likelihood L, the binarized prediction P and the ground truth T, these methods can make P to have more similar Betti numbers with the T. However, there is no guarantee that each persistent feature from L is matched to the correct one from T. This issue is prominent in grid-like data. For example, in a 2D image segmentation task, the ground truth mask is usually a binary image, which leads all points in $\mathcal{D}(T)$ to overlap at the upper-left corner of the diagram, i.e., (0, 1). As a result, the distances between each point in $\mathcal{D}(L)$ and any point in $\mathcal{D}(T)$ are the same, which causes confusions when matching the diagrams. We therefore propose to address this problem by introducing spatial awareness that further leverages the spatial distance between two persistent features in the original image when matching persistent diagrams.
@@ 3.2.1 Spatial-Aware Matching and Loss Formulation
Fig. 3 shows matching examples between T and L from a converged segmentation model (more detailed examples in supplementary material (SM)). From the limited examples that we manually compared, it seems the Betti-matching [31] is indeed a more accurate method. However, our proposed method still significantly reduces the matching error compared with the Wasserstein matching at a much smaller computational cost than the Betti-matching. Unfortunately, there is no ‘ground truth’ available to compare the matching accuracy numerically. Moreover, it is worth mentioning that although the matching on segmentation from a converged model offers a partial insight on this problem, it does not provide a comprehensive understanding of how the matching precisely influences the whole training process, which can be less straightforward to see. Accordingly, the final segmentation quality on testing images serves as the gold standard for the evaluation and empirically it is shown that a less confused matching enables better topological accuracy in the segmentation ([31], Table 1).
@@ 3.2.2 Gradient of SATLoss
where $\omega$ is the model weights, $L(c_{b}(\cdot))$ and $L(c_{d}(\cdot))$ are the pixel values of the creator and destroyer in L. In practice, the gradient pushes the value of creator pixel up and destroyer down if the persistent feature has a match in $\mathcal{D}(T)$ to prolong the persistence of the persistent feature. On contrary, if the persistent feature is matched to the diagonal, the gradient pushes the creator value down and the destroyer value up to shorten the persistence of this feature. The difference with Wasserstein matching is that our SATLoss enables the gradient to act more selectively on the persistent features and their critical cells, reducing the number of persistent features that are incorrectly prolonged or eliminated.
@@ 4.1 Datasets
Our main comparison with the SOTA topological segmentation loss functions excluding the BMLoss is performed over four diverse datasets and extends to more general tubular structure segmentation beyond biomedical images:
  • CREMI [9]: Segmentation of extracellular matrix of brain neurons in electronic microscopy images.
@@
Considering that the computational cost of persistent homology is still notably higher than conventional segmentation training objectives, it is a common practice in the literature of this field to crop the full image to small image patches for training and computing the evaluation metrics, for example, $64\times 64$ [29, 13]. However, very small patch size leads to limited semantic and topological information. In our experiments, we push it to the limit of our computational resources and use a patch size of around $300\times 300$ among training, validation and testing.
Table 2: Comparison with BMLoss on C.Elegan-small dataset.
[…]
@@
[…]
@@ 4.2 Implementation Details
Since the topic is the loss function, we follow [29, 13, 31, 12, 14] and adopt a consistent, general U-Net [28] network structure and set of hyperparameters for all experiments. An Adam optimizer is used with a weight decay of 1e-3 and an initial learning rate of 1e-3, decayed to 1e-4 at half of the total number of epochs. We train 30 epochs for Roads, 50 epochs for CREMI and CrackTree and 100 epochs for DRIVE. We use the best pixel loss reported in the corresponding paper for each of the baseline methods and use the BCE loss as pixel loss with our proposed method. The best weights $\lambda$ for the topological loss functions are determined by the ablation studies in Sec. 4.6. All experiments are repeated with 3 random seeds. For the smaller CrackTree and DRIVE datasets, we further adopt a 3-fold cross validation under each seed. In each independent random experiment, the model with the best Dice score on the validation set is used for testing and the average test scores and their standard deviations over these experiments are reported. Our implementation is under the PyTorch framework, where the computation of persistent homology relies on the GUDHI (Geometry Understanding in Higher Dimensions) library [26].
For C.Elegan-small and Roads-small datasets, we follow [31] to determine the network structures and training hyperparameters. We train all loss functions using the same hyperparameters reported in [31] without finetuning in our method. Additionally, in [31] experiments were run for only once. Consider the small dataset size, we instead use a 3-fold cross validation where each fold is repeated with 3 random seeds to make the results statistically more convincing.
@@ 4.3 Evaluation Metrics
Two types of metrics: volumetric and topological are used to evaluate the general segmentation quality and the topological accuracy, respectively. For volumetric metrics, we adopt the pixel-wise accuracy and the Dice score. We also use a clDice score [29], which although is a pixel-based metric, indirectly implies the topological correctness. For topological metrics, we use the 0-th and 1-st Betti number error ($\beta_{0}$, $\beta_{1}$), which are the most commonly-used measurements of topological accuracy. We stop using the Betti-matching metric proposed in [31] because the computational cost is too high given the comprehensive experiments.
@@ 4.5 Comparison with BMLoss
For comparison with BMLoss, we also include the clDice and TCLoss in the baseline, as the first method is the newest baseline in [31] and TCLoss is a newer method than BMLoss (which did not include BMLoss in their baseline). On another note, although the metrics are commonly computed on the image patches [13, 27, 29], in [31] they are computed over the full image. We therefore report the results on both the patch size P and full image size F. The results are shown in Table 3, Table 2 and Fig. 5. While our proposed method does not achieve the same performance level in global topological accuracy on full image size with BMLoss on the Roads-small dataset, on boarder evaluations which take the other results into consideration, our method demonstrates comparable, if not superior, performance. Furthermore, as reported in Table 4, our method achieves the results at a much smaller training time compared with BMLoss. Such a fundamental constraint of BMLoss is much less a problem of our method. It is important to note that in addition to the algorithm complexity, the coding language and implementation details can also influence the training time. The bottom-level codes in the GUDHI library we use to compute persistent homology are written in C++, while the codes for BMLoss are all written in Python. So this comparison is unfair in a strict sense. Given the complexity of the BMLoss, a reimplementation in C++ is challenging and is beyond the scope of this study. To the best of our knowledge, the Python version we used is the only and official implementation available. Therefore, despite the differences in implementation context, our proposed method still demonstrates superior computational efficiency.